# Solitary cluster waves in periodic potentials: Formation, propagation, and soliton-mediated particle transport

### v1

@@ VI.3 Soliton velocity
The corresponding expression for the type B mode of soliton motion is given in Appendix B.
An accurate treatment requires to consider the relaxation of the $n_{\rm b}$-clusters towards positions of stable mechanical equilibria. This will be discussed further below in Sec. IX in connection with an effective soliton-soliton interaction.
VII Number of solitons
@@
For determining $N_{\rm sol}^{\rm min}$, we consider the presoliton state with $\Delta=\Delta_{\rm min}\!-\!1$. It is composed of $N_{\rm b}(\Delta_{\rm min}\!-\!1)=N_{\rm pre}^{\rm b}$ clusters of the basic size $n_{\rm b}$ and $n_{\rm res}$ residual particles. By adding one particle, the presoliton state rearranges into a nonequilibrium steady state, where a maximal number of stable $n_{\rm b}$-clusters remains present. Differently speaking, this state carries a minimal number $N_{\rm sol}^{\rm min}$ of running solitons and an integer number $N_{\rm b}(\Delta_{\rm min})<N_{\rm b}(\Delta_{\rm min}\!-\!1)$ of clusters of size $n_{\rm b}$. As described in Sec. VI, all clusters involved in the soliton propagation are formed out of the basic and the core cluster, i.e. we need the particles of one basic stable cluster and the particles of one core cluster to generate one soliton. This means that
@@
For $\Delta_{\rm min}$ not satisfiying condition (33) or $\Delta_{\rm max}=\Delta_{\rm min}$, $\delta N_{\rm b}=\delta N_{\rm sol}=0$.
@@
VIII Magic particle sizes
@@
Figure 8 shows the results of this analysis. In addition, we have indicated by the color coding how many solitons form according to Eq. (53).
@@
IX Effective soliton-soliton interaction
@@
for a single soliton ($N_{\rm sol}=1$). The limit $L\to\infty$ corresponds to the approximation when disregarding the relaxation of the $n_{\rm b}$-clusters, i.e. $v_{\rm sol}(1,L\to\infty)=v_{\rm sol}^{\infty}=\lceil n_{\rm b}\sigma\rceil/\tau_{\rm sol}(y_{1},y_{2}^{\infty})$. For type B solitons, the expression for the relaxation-corrected velocity can be derived analogously, see Appendix B.
@@
X Soliton-mediated particle currents
@@
The derivation of $d$ is a bit more involved and given in Appendix C. The result is
@@
We have validated Eq. (60) for various parameter sets. As an example, we show in Fig. 11 simulated particle currents and $j_{\rm st}$ calculated from Eq. (60) as a function of $N_{\rm sol}$ for a system size $L=75$ and otherwise the same parameters as in Fig. 10. The number $N_{\rm sol}$ increases linearly with the overfilling $\Delta$ for $\Delta>\Delta_{\rm min}$, see the upper inset. The decrease of the single-soliton velocity $v_{\rm sol}$ with $N_{\rm sol}$ shown in the other inset is due to incomplete relaxation of $n_{\rm b}$-clusters. It leads to the sublinear increase of $j_{\rm st}$ with $N_{\rm sol}$ for $N_{\rm sol}>6$. The dashed line shows the behavior if the relaxation of $n_{\rm b}$-clusters is neglected. The slowing-down of $v_{\rm sol}$ with $N_{\rm sol}$ corresponds to the effective repulsive soliton-soliton interaction discussed in Sec. IX.
@@
XI Particle diameters larger than wavelength of potential
@@
A bit more care has to be taken when transforming soliton velocities and particle currents. For the sake of brevity, we refer to the system with the primed quantities as the “primed system”. The soliton velocity in Eq. (40) without relaxation correction becomes $v^{\prime}_{\rm sol}=v_{\rm sol}+n_{\rm b}\lfloor\sigma^{\prime}\rfloor/\tau^{\prime}_{\rm sol}$, where the time period of the soliton motion does not change, $\tau^{\prime}_{\rm sol}=\tau_{\rm sol}$. The relaxation correction for the soliton velocity in the primed system can be determined analogously to that in the unprimed system as described in Sec. IX. In Eq. (56), one needs to replace $L$ and $\sigma$ by $L^{\prime}$ and $\sigma^{\prime}$, and can take $y_{1}^{\prime}=y_{1}$. It then becomes a determining for $y_{2}^{\prime}(L^{\prime})$. Replacing all unprimed by the primed quantities in Eqs. (37) and (57) gives the relaxation-corrected single soliton velocity $v^{\prime}_{\rm sol}(1,L^{\prime})=\lceil n_{\rm b}\sigma^{\prime}\rceil/\tau_{\rm sol}(y_{1},y_{2}^{\prime}(L^{\prime}))$. Using this in Eq. (58), one obtains $v_{\rm sol}^{\prime}(N_{\rm sol},L^{\prime})=v_{\rm sol}^{\prime}(1,L^{\prime}/N_{\rm sol})$. As for the particle current in Eq. (60), $d$ and $k$ as well as $N$ and $N_{\rm sol}$ are unchanged, i.e. one needs to replace $L$ by $L^{\prime}$ and $v_{\rm sol}(N_{\rm sol},L)$ by $v_{\rm sol}^{\prime}(N_{\rm sol},L^{\prime})$.
XII Conclusions
@@
The presoliton state is formed by evenly spaced mechanically stable $n_{\rm b}$-clusters, i.e. clusters formed by $n_{\rm b}$ particles in contact. Surprisingly, mechanical stabilizability of a cluster is related to a purely geometric property, namely the residual free space left in potential wells accommodating the cluster. As a consequence, the number $n_{\rm b}$ is determined by a principle of minimum free residual space. Using this principle, we derived an explicit expression for $n_{\rm b}$.
We believe that the connection between mechanical stabilizability and free residual space is valid for all symmetric potentials with a single period. Generally, one may ask whether laws exist for largest mechanically stable packings of hard spheres in periodic potentials. This question can be viewed as an extension of the well-known closest packing problem of hard spheres in free space.
In the running state, the cluster sequence in the soliton propagation has a smallest core cluster made of $n_{\rm c}$ particles. All other clusters involved are composed of one $n_{\rm c}$-cluster and $n_{\rm b}$-clusters. The size $n_{\rm c}$ follows from the requirement of barrier-free motions of soliton clusters. The simplest soliton propagation mode is that of alternating movements of an $n_{\rm c}$- and a composite $(n_{\rm c}+n_{\rm b})$-cluster. It resembles the situation in a relay race.
@@
The incomplete relaxation of $n_{\rm b}$-clusters towards their positions of mechanical equilibria in a running state gives rise to an effective repulsive soliton-soliton interaction. This effect, recently observed experimentally [21], is also responsible for a slowing down of particle currents.
@@
We expect the theoretical concepts presented here to be relevant also for particle transport under time-dependent driving and in dimensions higher than one. For the numerous examples given in the Introduction [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], they can trigger new ways for analyzing cluster dynamics on a microscopic level.
Acknowledgements.
Financial support by the Czech Science Foundation (Project No. 23-09074L) and the Deutsche Forschungsgemeinschaft (Project No. 521001072) is gratefully acknowledged.
@@ Appendix B Velocity of type B solitons
For deriving $d$, we need to specify positions of the soliton. As initial particle configuration we consider one, where the soliton $n_{\rm c}$-cluster attaches to an $n_{\rm b}$-cluster, see the first line in Figs. C.1(a) and (b). In the respective $n_{\rm b}$-cluster, we tag a particle and calculate how it is displaced after successive soliton circulations. Each circulation is finished when the soliton $n_{\rm c}$-cluster attaches to the $n_{\rm b}$-cluster that contains the tagged particle. Summing up all displacements of the tagged particle in the $k$ soliton circulations, with $k$ from Eq. (64), we obtain $d$.
Let $s$ be the order number of the tagged particle and $y$ the position of the soliton ($n_{\rm c}$-cluster) after the $j$th soliton circulation, $j=0,\ldots,k-1$, see Fig. C.1. The position of the tagged particle in this configuration is
@@
Figure C.1: Simulated particle configurations for the same parameters $\sigma=0.57$, $f=0.01$ and $L=40$ as in Fig. 5(a) to illustrate the derivation of the soliton-mediated particle current ($N_{\rm sol}=1$, $n_{\rm b}=5$, $n_{\rm c}=2$). In the first line, the $n_{\rm c}$-cluster attaches to an $n_{\rm b}$-cluster, where we have tagged the first particle (order number $s=0$, orange) and the fourth particle ($s=3$, blue). The second and third lines show the configurations after one soliton circulation for the orange and blue tagged particles, respectively. After the soliton circulation, the orange tagged particle has order number $s^{\prime}=s+n_{\rm c}$ [change $\Delta s=n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}=2$, Eq. (62)], and the blue tagged particle has order number $s^{\prime}=0$ [change $\Delta s=n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}-n_{\rm b}=n_{\rm c}-n_{\rm b}=-3$, Eq. (63)]. Vertical dotted lines indicate the displacements $\Delta x_{1}$ and $\Delta x_{2}$ of the tagged particles corresponding to their different type of order number change, see Eq. (85).

### v2

@@ 6.3 Soliton velocity
The corresponding expression for the type B mode of soliton motion is given in B.
An accurate treatment requires to consider the relaxation of the $n_{\rm b}$-clusters towards positions of stable mechanical equilibria. This will be discussed further below in Sec. 9 in connection with an effective soliton-soliton interaction.
7 Number of solitons
@@
For determining $N_{\rm sol}^{\rm min}$, we consider the presoliton state with $\Delta=\Delta_{\rm min}\!-\!1$. It is composed of $N_{\rm b}(\Delta_{\rm min}\!-\!1)=N_{\rm pre}^{\rm b}$ clusters of the basic size $n_{\rm b}$ and $n_{\rm res}$ residual particles. By adding one particle, the presoliton state rearranges into a nonequilibrium steady state, where a maximal number of stable $n_{\rm b}$-clusters remains present. Differently speaking, this state carries a minimal number $N_{\rm sol}^{\rm min}$ of running solitons and an integer number $N_{\rm b}(\Delta_{\rm min})<N_{\rm b}(\Delta_{\rm min}\!-\!1)$ of clusters of size $n_{\rm b}$. As described in Sec. 6, all clusters involved in the soliton propagation are formed out of the basic and the core cluster, i.e. we need the particles of one basic stable cluster and the particles of one core cluster to generate one soliton. This means that
@@
For $\Delta_{\rm min}$ not satisfying condition (33) or $\Delta_{\rm max}=\Delta_{\rm min}$, $\delta N_{\rm b}=\delta N_{\rm sol}=0$.
@@
8 Magic particle sizes
@@
Figure 8 shows the results of this analysis. In addition, we have indicated by the color coding how many solitons form according to Eq. (53). We suggest to test these predictions in experiments.
@@
9 Effective soliton-soliton interaction
@@
for a single soliton ($N_{\rm sol}=1$). The limit $L\to\infty$ corresponds to the approximation when disregarding the relaxation of the $n_{\rm b}$-clusters, i.e. $v_{\rm sol}(1,L\to\infty)=v_{\rm sol}^{\infty}=\lceil n_{\rm b}\sigma\rceil/\tau_{\rm sol}(y_{1},y_{2}^{\infty})$. For type B solitons, the expression for the relaxation-corrected velocity can be derived analogously, see B.
@@
10 Soliton-mediated particle currents
@@
The derivation of $d$ is a bit more involved and given in C. The result is
@@
We have validated Eq. (60) for various parameter sets. As an example, we show in Fig. 11 simulated particle currents and $j_{\rm st}$ calculated from Eq. (60) as a function of $N_{\rm sol}$ for a system size $L=75$ and otherwise the same parameters as in Fig. 10. The number $N_{\rm sol}$ increases linearly with the overfilling $\Delta$ for $\Delta>\Delta_{\rm min}$, see the upper inset. The decrease of the single-soliton velocity $v_{\rm sol}$ with $N_{\rm sol}$ shown in the other inset is due to incomplete relaxation of $n_{\rm b}$-clusters. It leads to the sublinear increase of $j_{\rm st}$ with $N_{\rm sol}$ for $N_{\rm sol}>6$. The dashed line shows the behavior if the relaxation of $n_{\rm b}$-clusters is neglected. The slowing-down of $v_{\rm sol}$ with $N_{\rm sol}$ corresponds to the effective repulsive soliton-soliton interaction discussed in Sec. 9.
@@
The main results of this section are Eq. (60) for the soliton-mediated current in finite systems and Eq. (69) for the current-density relation in the thermodynamic limit. The latter predicts a dynamical phase transition to occur from a jammed to a current-carrying state at a particle number density $\rho_{\rm b}=n_{\rm b}/\lceil n_{\rm b}\sigma\rceil$ for $k_{\rm B}T/U_{0}\ll 1$, strictly speaking in the limit $k_{\rm B}T/U_{0}\to 0$. We believe that the strong increase of currents found in earlier investigations of current-density relations at $k_{\rm B}T/U_{0}=1/6$ for $\rho>1$ [47] is reflecting this transition. It would be interesting to see whether it can be obtained by continuum theories, as, e.g., the dynamical density functional theory [48], the power density functional theory [49, 50], or the macroscopic fluctuation theory [51].
11 Particle diameters larger than wavelength of potential
@@
A bit more care has to be taken when transforming soliton velocities and particle currents. For the sake of brevity, we refer to the system with the primed quantities as the “primed system”. The soliton velocity in Eq. (40) without relaxation correction becomes $v^{\prime}_{\rm sol}=v_{\rm sol}+n_{\rm b}\lfloor\sigma^{\prime}\rfloor/\tau^{\prime}_{\rm sol}$, where the time period of the soliton motion does not change, $\tau^{\prime}_{\rm sol}=\tau_{\rm sol}$. The relaxation correction for the soliton velocity in the primed system can be determined analogously to that in the unprimed system as described in Sec. 9. In Eq. (56), one needs to replace $L$ and $\sigma$ by $L^{\prime}$ and $\sigma^{\prime}$, and can take $y_{1}^{\prime}=y_{1}$. It then becomes a determining for $y_{2}^{\prime}(L^{\prime})$. Replacing all unprimed by the primed quantities in Eqs. (37) and (57) gives the relaxation-corrected single soliton velocity $v^{\prime}_{\rm sol}(1,L^{\prime})=\lceil n_{\rm b}\sigma^{\prime}\rceil/\tau_{\rm sol}(y_{1},y_{2}^{\prime}(L^{\prime}))$. Using this in Eq. (58), one obtains $v_{\rm sol}^{\prime}(N_{\rm sol},L^{\prime})=v_{\rm sol}^{\prime}(1,L^{\prime}/N_{\rm sol})$. As for the particle current in Eq. (60), $d$ and $k$ as well as $N$ and $N_{\rm sol}$ are unchanged, i.e. one needs to replace $L$ by $L^{\prime}$ and $v_{\rm sol}(N_{\rm sol},L)$ by $v_{\rm sol}^{\prime}(N_{\rm sol},L^{\prime})$.
12 Conclusions
@@
The presoliton state is formed by evenly spaced mechanically stable $n_{\rm b}$-clusters, i.e. clusters formed by $n_{\rm b}$ particles in contact. Surprisingly, mechanical stabilizability of a cluster is related to a purely geometric property, namely the residual free space left in potential wells accommodating the cluster. As a consequence, the number $n_{\rm b}$ is determined by a principle of minimum free residual space. Using this principle, we derived an explicit expression for $n_{\rm b}$, see Eqs. (23), (26).
We believe that there exists a connection between mechanical stabilizability and free residual space for all symmetric potentials with a single period, maybe with a generalization that the presoliton state is formed by more than one basic cluster. Generally, one may ask whether laws exist for largest mechanically stable packings of hard spheres in periodic potentials. This question can be viewed as an extension of the well-known closest packing problem of hard spheres in free space.
In the running state, the cluster sequence in the soliton propagation has a smallest core cluster made of $n_{\rm c}$ particles. All other clusters involved are composed of one $n_{\rm c}$-cluster and $n_{\rm b}$-clusters. The size $n_{\rm c}$ follows from the requirement of barrier-free motions of soliton clusters, which for infinitesimal weak forces is given by Eq. (34). The simplest soliton propagation mode is that of alternating movements of an $n_{\rm c}$- and a composite $(n_{\rm c}+n_{\rm b})$-cluster. It resembles the situation in a relay race: the $n_{\rm c}$-cluster can by viewed as the relay that is carried by a core runner and passed to another runner, when an $n_{\rm c}$-cluster attaches to the left end of an $n_{\rm b}$-cluster and an $(n_{\rm c}+n_{\rm b})$-cluster forms. Subsequently the relay is passed to another core runner, when the $n_{\rm c}$-cluster detaches from the right end of the $(n_{\rm c}+n_{\rm b})$-cluster, and so on.
@@
The incomplete relaxation of $n_{\rm b}$-clusters towards their positions of mechanical equilibria in a running state gives rise to an effective repulsive soliton-soliton interaction, which corresponds to a decrease of soliton velocities as described by Eqs. (57) [or (81)] and (58). This effect, recently observed experimentally [22], is also responsible for a slowing down of particle currents, see Fig. 11 and Eqs. (60), (69).
@@
We expect the theoretical concepts presented here to be relevant also for particle transport under time-dependent driving and in dimensions higher than one. For the numerous examples given in the Introduction [1, 2, 3, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], as well as in new setups of twisted optical and magnetic lattices [56, 57], they can trigger new ways for analyzing cluster dynamics on a microscopic level.
Acknowledgements
Financial support by the Czech Science Foundation (Project No. 23-09074L) and the Deutsche Forschungsgemeinschaft (Project No. 521001072) is gratefully acknowledged. We thank P. Tierno for very helpful discussions on the impact of our results for experiments.
@@ Appendix B Velocity of type B solitons
Figure B.1: Simulated particle configurations for the same parameters $\sigma=0.57$, $f=0.01$ and $L=40$ as in Fig. 5(a) to illustrate the derivation of the soliton-mediated particle current ($N_{\rm sol}=1$, $n_{\rm b}=5$, $n_{\rm c}=2$). In the first line, the $n_{\rm c}$-cluster attaches to an $n_{\rm b}$-cluster, where we have tagged the first particle (order number $s=0$, orange) and the fourth particle ($s=3$, blue). The second and third lines show the configurations after one soliton circulation for the orange and blue tagged particles, respectively. After the soliton circulation, the orange tagged particle has order number $s^{\prime}=s+n_{\rm c}$ [change $\Delta s=n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}=2$, Eq. (62)], and the blue tagged particle has order number $s^{\prime}=0$ [change $\Delta s=n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}-n_{\rm b}=n_{\rm c}-n_{\rm b}=-3$, Eq. (63)]. Vertical dotted lines indicate the displacements $\Delta x_{1}$ and $\Delta x_{2}$ of the tagged particles corresponding to their different type of order number change, see Eq. (85).
@@
For deriving $d$, we need to specify positions of the soliton. As initial particle configuration we consider one, where the soliton $n_{\rm c}$-cluster attaches to an $n_{\rm b}$-cluster, see the first line in Figs. B.1(a) and (b). In the respective $n_{\rm b}$-cluster, we tag a particle and calculate how it is displaced after successive soliton circulations. Each circulation is finished when the soliton $n_{\rm c}$-cluster attaches to the $n_{\rm b}$-cluster that contains the tagged particle. Summing up all displacements of the tagged particle in the $k$ soliton circulations, with $k$ from Eq. (64), we obtain $d$.
Let $s$ be the order number of the tagged particle and $y$ the position of the soliton ($n_{\rm c}$-cluster) after the $j$th soliton circulation, $j=0,\ldots,k-1$, see Fig. B.1. The position of the tagged particle in this configuration is